# Dynamical Edge Modes in $p$-form Gauge Theories

### v1

@@ Abstract
We extend our recently identified dynamical edge mode boundary condition to $p$-form gauge theories, revealing their edge modes as Goldstone bosons arising from gauge transformations with support on the boundary. The symplectic conjugates of these edge modes correspond to the electric-field-like components normal to the boundary. We demonstrate that both the symplectic form and the Hamiltonian naturally decompose into bulk and edge parts. When the boundary is a stretched horizon, we show that the thermal edge partition function reduces to that of a codimension-two ghost $(p-1)$-form residing on the bifurcation surface. These findings provide a dynamical framework that elucidates observations made by several authors. Additionally, we generalize Donnelly and Wall’s non-dynamical approach to obtain edge partition functions for both massive and massless $p$-forms. In the context of a de Sitter static patch, these results are consistent with the edge partition functions found by several authors in arbitrary dimensions.
@@ 1 Introduction
In our previous work [4], we presented a dynamical framework for describing edge modes in Maxwell theory in the presence of a timelike boundary. Central to this framework is the introduction of what we call the dynamical edge mode (DEM) boundary condition, allowing for physical large gauge transformations and their symplectic conjugates. The goal of this paper is to extend this analysis to $p$-form gauge theories for any $p\geq 1$. These theories are intriguing due to their rich geometric structure [34] and electromagnetic duality, as well as their pivotal roles in various other areas of physics. For instance, the low-energy effective descriptions of string theory and M-theory—namely, 10D and 11D supergravity—feature several massless $p$-form gauge fields. In fact, the relevance of $p$-forms to entanglement entropy in string theory was part of the motivation for this work [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. Additionally, $p$-form gauge theories provide quintessential examples of higher-form symmetries [47]. And in even spacetime dimensions $D$, a $p$-form gauge theory with $p=\frac{D-2}{2}$ is conformally invariant, thereby providing an infinite family of CFTs.
Despite their potential roles in the aforementioned topics, edge modes in the case of $p>1$ have not been as thoroughly studied as their Maxwell ($p=1$) counterpart. Nonetheless, some calculations of partition functions and entanglement entropy [48, 49, 50, 2, 14, 51, 52] suggest the existence of edge contributions from a $(p-1)$-form residing on the codimension-2 entangling surface, naturally generalizing the $p=1$ case. As we shall see, a natural extension of our DEM framework in [4] nicely captures the edge contributions, thus providing a dynamical understanding of these results.
@@ 1.1 Overview of paper
splits cleanly into bulk and edge parts. Specifically, in section 3.1 we study the phase space on a Cauchy slice $\Sigma$, and we show that the degrees of freedom can be decomposed into bulk modes realized as $p$-forms living on $\Sigma$ and edge modes parametrized by $(p-1)$-forms $\alpha$, $E_{\perp}$ living on $\partial\Sigma$. Here $\alpha$ shifts under large gauge transformations and $E_{\perp}$ is the “electric” flux through the boundary. The symplectic form splits under this decomposition,
@@
Here $K$ is a certain pseudo-differential operator on $\partial\Sigma$, defined in (3.35). In section 3.3 we study the horizon limit, in which the boundary $\partial M$ approaches a bifurcate Killing horizon, and we find that $K$ simplifies to a multiple of the $(p-1)$-form Laplacian on $\partial\Sigma$,
@@
With these ingredients in hand, in section 4 we quantize and study thermal partition functions. The partition function manifestly factorizes into bulk and edge pieces, and in section 4.1 we evaluate the edge part, finding in (4.20) that it amounts to the reciprocal of a codimension-two $(p-1)$-form partition function on $\partial\Sigma$:
@@ 2 Review of $p$-form gauge theory
In this section we review $p$-form gauge theory and its quantization on a $D$-dimensional manifold. We write $A$ for the $p$-form gauge field and $F=dA$ for its $(p+1)$-form field strength. We take the gauge group to be $U(1)$ with fundamental charge $q$. We assume $p\leq D-2$ for simplicity. [^2] The case of $p=D-1$ is qualitatively different; naïvely it would be dual to a $\tilde{p}$-form gauge theory with $\tilde{p}=D-2-p=-1$. A well-known example is Maxwell in $D=2$, where the theory depends only on the volume of the manifold, not its specific geometry. This quasi-topological nature extends to cases where $p=D-1$, which make them extremely tractable and a popular starting point for testing new ideas in gauge theory [17, 20].
@@ 2.2 Number of degrees of freedom
It will be useful later as a sanity check to know the number of local degrees of freedom, or polarizations, of a $p$-form gauge field. It can be deduced as follows. A $p$-form initially has (${D\choose p}$ independent components but, at least locally, we can use gauge transformations to kill the time component, $t^{\mu}A_{\mu\dots}=0$. This is called temporal gauge. We still have residual time-independent gauge transformations, and we can use them to further set $v^{\mu}A_{\mu\dots}=0$ on shell, for some fixed spatial vector field $v^{\mu}$, perhaps corresponding to a particular coordinate. This leaves us with
@@
independent physical polarizations. Note that $p=1$ gives (${D-2\choose 1}=D-2$ polarizations, which is correct. This quick argument informs us about the number of $D$-dimensional degrees of freedom, but says nothing about the edge modes, which are effectively lower-dimensional degrees of freedom.
@@ 2.3 Euclidean partition function
$\mathcal{Z}_{\rm PI}^{p=1}[\mathcal{M}]=\sum_{\mathscr{F}\in\frac{2\pi}{q}\mathcal{H}^{2}(\mathcal{M},\mathbb{Z})}e^{-S[\mathscr{F}]}\det^{\prime}\left(\Delta_{0}\right)^{\frac{1}{2}}\left|\frac{H^{1}(\mathcal{M},\mathbb{R})}{\frac{2\pi}{q}H^{1}(\mathcal{M},\mathbb{Z})}\right|\frac{1}{|U(1)|}\int\mathcal{D}A_{\rm co}\,e^{-S[dA_{\rm co}]}\;.$ (2.21)
@@
$\mathcal{Z}_{\rm PI}^{p=1}[\mathcal{M}]=\frac{\det^{\prime}\left(\Delta_{0}\right)^{\frac{1}{2}}}{\det^{\prime}\left(\Delta_{1}^{T}\right)^{\frac{1}{2}}}\,\frac{1}{|U(1)|}\left|\frac{H^{1}(\mathcal{M},\mathbb{R})}{\frac{2\pi}{q}H^{1}(\mathcal{M},\mathbb{Z})}\right|\sum_{\mathscr{F}\in\frac{2\pi}{q}\mathcal{H}^{2}(\mathcal{M},\mathbb{Z})}e^{-S[\mathscr{F}]}\;.$ (2.23)
@@
while the integration of the exact part of a $k$-form naturally cancels against the co-exact part of a $(k-1)$-form up to a Jacobian factor
$\frac{\int\mathcal{D}A_{\rm ex}}{\int\mathcal{D}\lambda_{\rm co}}=\det^{\prime}\left(\Delta_{1}^{T}\right)^{\frac{1}{2}}\;,\qquad\frac{\int\mathcal{D}\lambda_{\rm ex}}{\int\mathcal{D}\phi_{\rm co}}=\det^{\prime}\left(\Delta_{0}\right)^{\frac{1}{2}}\;.$ (2.28)
@@
$=\sum_{\mathscr{F}\in\frac{2\pi}{q}\mathcal{H}^{3}(\mathcal{M},\mathbb{Z})}e^{-S[\mathscr{F}]}\det^{\prime}\left(\Delta_{1}^{T}\right)^{\frac{1}{2}}\left|\frac{H^{2}(\mathcal{M},\mathbb{R})}{\frac{2\pi}{q}H^{2}(\mathcal{M},\mathbb{Z})}\right|\det^{\prime}\left(\Delta_{0}\right)^{-\frac{1}{2}}\left|\frac{H^{1}(\mathcal{M},\mathbb{R})}{\frac{2\pi}{q}H^{1}(\mathcal{M},\mathbb{Z})}\right|^{-1}|U(1)|\int\mathcal{D}A_{\rm co}\,e^{-S[dA_{\rm co}]}\;.$ (2.29)
@@
$=\sum_{\mathscr{F}\in\frac{2\pi}{q}\mathcal{H}^{3}(\mathcal{M},\mathbb{Z})}e^{-S[\mathscr{F}]}\det^{\prime}\left(\Delta_{2}^{T}\right)^{-\frac{1}{2}}\left|\frac{H^{2}(\mathcal{M},\mathbb{R})}{\frac{2\pi}{q}H^{2}(\mathcal{M},\mathbb{Z})}\right|\det^{\prime}\left(\Delta_{1}^{T}\right)^{\frac{1}{2}}\left|\frac{H^{1}(\mathcal{M},\mathbb{R})}{\frac{2\pi}{q}H^{1}(\mathcal{M},\mathbb{Z})}\right|^{-1}\det^{\prime}\left(\Delta_{0}\right)^{-\frac{1}{2}}|U(1)|\;.$ (2.30)
@@
$\mathcal{Z}_{\rm PI}^{p}[\mathcal{M}]$ $=\sum_{\mathscr{F}\in\frac{2\pi}{q}\mathcal{H}^{p+1}(\mathcal{M},\mathbb{Z})}e^{-S[\mathscr{F}]}\int\frac{\mathcal{D}A}{|\mathcal{G}_{p-1}/(\mathcal{G}_{p-2}/\dots)|}\,e^{-S[dA]}$ (2.31)
$=\sum_{\mathscr{F}\in\frac{2\pi}{q}\mathcal{H}^{p+1}(\mathcal{M},\mathbb{Z})}e^{-S[\mathscr{F}]}\prod_{k=0}^{p}\left(\det^{\prime}\left(\Delta_{k}^{T}\right)^{-\frac{1}{2}}\left|\frac{H^{k}(\mathcal{M},\mathbb{R})}{\frac{2\pi}{q}H^{k}(\mathcal{M},\mathbb{Z})}\right|\right)^{(-)^{p-k}}\;.$
@@
$=|\mathcal{G}_{\rm nontriv}|\left|\frac{H^{p-1}(\mathcal{M},\mathbb{R})}{\frac{2\pi}{q}H^{p-1}(\mathcal{M},\mathbb{Z})}\right|\,\prod_{k=0}^{p-2}\left(\det^{\prime}\left(\Delta_{k}^{T}\right)^{-\frac{1}{2}}\left|\frac{H^{k}(\mathcal{M},\mathbb{R})}{\frac{2\pi}{q}H^{k}(\mathcal{M},\mathbb{Z})}\right|\right)^{(-)^{p-k+1}}$
where $\mathcal{G}_{\rm nontriv}$ is the nontrivial part of $\mathcal{G}_{p-1}$, i.e. $\frac{2\pi}{q}\mathcal{H}^{p}(\mathcal{M},\mathbb{Z})$ plus $d\lambda$ with $\lambda$ a non-closed $(p-1)$-form. Then the measure can be written simply as $\mathcal{D}A/|\mathcal{G}|$. For later use we evaluate $|U(1)|$ explicitly here. A natural measure for the scalar gauge parameter $\phi$ is $\mathcal{D}\phi=\prod_{n}\frac{\mu^{p+1}d\phi_{n}}{\sqrt{2\pi}}$, associated with the Laplacian eigendecomposition $\phi(x)=\sum_{n}\phi_{n}f_{n}(x)$. Here the arbitrary mass scale $\mu$ is analogous to a UV cutoff, and is included so that the measure is dimensionless. The normalized zero mode is $f_{0}(x)=V_{\mathcal{M}}^{-1/2}$, where $V_{\mathcal{M}}$ is the volume of $\mathcal{M}$. The zero mode coefficient then inherits the periodicity $\phi_{0}\sim\phi_{0}+\frac{2\pi}{q}V_{\mathcal{M}}^{1/2}$. We then have
[…]
@@
$\det^{\prime}\left(\Delta_{k}\right)=\det^{\prime}\left(\Delta_{k}^{T}\right)\det^{\prime}\left(\Delta_{k-1}^{T}\right)\;.$ (2.34)
The functional determinant factors tell us about the local degrees of freedom of the theory. We now carry out a sanity check to confirm that in total there are (${D-2\choose p}$ local degrees of freedom, as argued above. A $k$-form has (${D\choose k}$ independent components, so a determinant $\det^{\prime}\left(\Delta_{k}\right)^{-\frac{1}{2}}$ represents (${D\choose k}$ degrees of freedom. Similarly its reciprocal represents ($-{D\choose k}$ degrees of freedom. Let us write the number of degrees of freedom in a transverse $k$-form determinant as $n_{k}$. From (2.34) we apparently have $n_{k}+n_{k-1}={D\choose k}$. This recursion relation is solved by

### v2

@@ Abstract
We extend our recently identified dynamical edge mode boundary condition to $p$-form gauge theories, revealing their edge modes as Goldstone bosons arising from gauge transformations with support on the boundary. The symplectic conjugates of these edge modes correspond to the electric-field-like components normal to the boundary. We demonstrate that both the symplectic form and the Hamiltonian naturally decompose into bulk and edge parts. When the boundary is a stretched horizon, we show that the thermal edge partition function reduces to that of a codimension-two ghost (p1)(p-1)-form residing on the bifurcation surface. These findings provide a dynamical framework that elucidates observations made by several authors. Additionally, we generalize Donnelly and Wall’s non-dynamical approach to obtain edge partition functions for both massive and massless $p$-forms. In the context of a de Sitter static patch, these results are consistent with the edge partition functions found by several authors in arbitrary dimensions.
@@ 1 Introduction
In our previous work [4], we presented a dynamical framework for describing edge modes in Maxwell theory in the presence of a timelike boundary. Central to this framework is the introduction of what we call the dynamical edge mode (DEM) boundary condition, allowing for physical large gauge transformations and their symplectic conjugates. The goal of this paper is to extend this analysis to $p$-form gauge theories for any $p\geq 1$. These theories are intriguing due to their rich geometric structure [34] and electromagnetic duality, as well as their pivotal roles in various other areas of physics. For instance, the low-energy effective descriptions of string theory and M-theory—namely, 10D and 11D supergravity—feature several massless $p$-form gauge fields. In fact, the relevance of $p$-forms to entanglement entropy in string theory was part of the motivation for this work [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. Additionally, $p$-form gauge theories provide quintessential examples of higher-form symmetries [47]. And in even spacetime dimensions $D$, a $p$-form gauge theory with $p=\frac{D-2}{2}$ is conformally invariant, thereby providing an infinite family of CFTs. It is also worth noting that edge modes have been studied in $p$-form BF theories [48, 49].
Despite their potential roles in the aforementioned topics, edge modes in the case of $p>1$ have not been as thoroughly studied as their Maxwell ($p=1$) counterpart. Nonetheless, some calculations of partition functions and entanglement entropy [50, 51, 52, 2, 14, 53, 54] suggest the existence of edge contributions from a (p1)(p-1)-form residing on the codimension-2 entangling surface, naturally generalizing the $p=1$ case. As we shall see, a natural extension of our DEM framework in [4] nicely captures the edge contributions, thus providing a dynamical understanding of these results.
@@ 1.1 Overview of paper
splits cleanly into bulk and edge parts. Specifically, in section 3.1 we study the phase space on a Cauchy slice $\Sigma$, and we show that the degrees of freedom can be decomposed into bulk modes realized as $p$-forms living on $\Sigma$ and edge modes parametrized by (p1)(p-1)-forms $\alpha$, $E_{\perp}$ living on $\partial\Sigma$. Here $\alpha$ shifts under large gauge transformations and $E_{\perp}$ is the “electric” flux through the boundary. The symplectic form splits under this decomposition,
@@
Here $K$ is a certain pseudo-differential operator on $\partial\Sigma$, defined in (3.35). In section 3.3 we study the horizon limit, in which the boundary $\partial M$ approaches a bifurcate Killing horizon, and we find that $K$ simplifies to a multiple of the (p1)(p-1)-form Laplacian on $\partial\Sigma$,
@@
With these ingredients in hand, in section 4 we quantize and study thermal partition functions. The partition function manifestly factorizes into bulk and edge pieces, and in section 4.1 we evaluate the edge part, finding in (4.20) that it amounts to the reciprocal of a codimension-two (p1)(p-1)-form partition function on $\partial\Sigma$:
@@ 2 Review of $p$-form gauge theory
In this section we review $p$-form gauge theory and its quantization on a $D$-dimensional manifold. We write $A$ for the $p$-form gauge field and $F=dA$ for its $(p+1)$-form field strength. We take the gauge group to be $U(1)$ with fundamental charge $q$. We assume $p\leq D-2$ for simplicity. [^2] The case of p=D1p=D-1 is qualitatively different; naïvely it would be dual to a $\tilde{p}-$form gauge theory with $\tilde{p}=D-2-p=-1$. A well-known example is Maxwell in $D=2$, where the theory depends only on the volume of the manifold, not its specific geometry. This quasi-topological nature extends to cases where p=D1p=D-1, which make them extremely tractable and a popular starting point for testing new ideas in gauge theory [55, 17, 20].
@@ 2.2 Number of degrees of freedom
It will be useful later as a sanity check to know the number of local degrees of freedom, or polarizations, of a $p$-form gauge field. It can be deduced as follows. A $p$-form initially has $D\choose p$} independent components but, at least locally, we can use gauge transformations to kill the time component, $t^{\mu}A_{\mu\dots}=0$. This is called temporal gauge. We still have residual time-independent gauge transformations, and we can use them to further set $v^{\mu}A_{\mu\dots}=0$ on shell, for some fixed spatial vector field $v^{\mu}$, perhaps corresponding to a particular coordinate. This leaves us with
@@
independent physical polarizations. Note that $p=1$ gives ${D-2\choose 1}=D-2$ polarizations, which is correct. This quick argument informs us about the number of $D$-dimensional degrees of freedom, but says nothing about the edge modes, which are effectively lower-dimensional degrees of freedom.
@@ 2.3 Euclidean partition function
$\mathcal{Z}_{\rm PI}^{p=1}[\mathcal{M}]=\sum_{\mathscr{F}\in\frac{2\pi}{q}\mathcal{H}^{2}(\mathcal{M},\mathbb{Z})}e^{-S[\mathscr{F}]}\det\hskip 0.0pt^{\prime}\left(\Delta_{0}\right)^{\frac{1}{2}}\left|\frac{H^{1}(\mathcal{M},\mathbb{R})}{\frac{2\pi}{q}H^{1}(\mathcal{M},\mathbb{Z})}\right|\frac{1}{|U(1)|}\int\mathcal{D}A_{\rm co}\,e^{-S[dA_{\rm co}]}\;.$ (2.21)
@@
$\mathcal{Z}_{\rm PI}^{p=1}[\mathcal{M}]=\frac{\det\hskip 0.0pt^{\prime}\left(\Delta_{0}\right)^{\frac{1}{2}}}{\det\hskip 0.0pt^{\prime}\left(\Delta_{1}^{T}\right)^{\frac{1}{2}}}\,\frac{1}{|U(1)|}\left|\frac{H^{1}(\mathcal{M},\mathbb{R})}{\frac{2\pi}{q}H^{1}(\mathcal{M},\mathbb{Z})}\right|\sum_{\mathscr{F}\in\frac{2\pi}{q}\mathcal{H}^{2}(\mathcal{M},\mathbb{Z})}e^{-S[\mathscr{F}]}\;.$ (2.23)
@@
while the integration of the exact part of a $k$-form naturally cancels against the co-exact part of a (k1)(k-1)-form up to a Jacobian factor
$\frac{\int\mathcal{D}A_{\rm ex}}{\int\mathcal{D}\lambda_{\rm co}}=\det\hskip 0.0pt^{\prime}\left(\Delta_{1}^{T}\right)^{\frac{1}{2}}\;,\qquad\frac{\int\mathcal{D}\lambda_{\rm ex}}{\int\mathcal{D}\phi_{\rm co}}=\det\hskip 0.0pt^{\prime}\left(\Delta_{0}\right)^{\frac{1}{2}}\;.$ (2.28)
@@
$=\hskip-5.69054pt\sum_{\mathscr{F}\in\frac{2\pi}{q}\mathcal{H}^{3}(\mathcal{M},\mathbb{Z})}\hskip-14.22636pte^{-S[\mathscr{F}]}\det\hskip 0.0pt^{\prime}\left(\Delta_{1}^{T}\right)^{\frac{1}{2}}\left|\frac{H^{2}(\mathcal{M},\mathbb{R})}{\frac{2\pi}{q}H^{2}(\mathcal{M},\mathbb{Z})}\right|\det\hskip 0.0pt^{\prime}\left(\Delta_{0}\right)^{-\frac{1}{2}}\left|\frac{H^{1}(\mathcal{M},\mathbb{R})}{\frac{2\pi}{q}H^{1}(\mathcal{M},\mathbb{Z})}\right|^{-1}|U(1)|\int\mathcal{D}A_{\rm co}\,e^{-S[dA_{\rm co}]}\;.$ (2.29)
@@
$=\hskip-5.69054pt\sum_{\mathscr{F}\in\frac{2\pi}{q}\mathcal{H}^{3}(\mathcal{M},\mathbb{Z})}\hskip-14.22636pte^{-S[\mathscr{F}]}\det\hskip 0.0pt^{\prime}\left(\Delta_{2}^{T}\right)^{-\frac{1}{2}}\left|\frac{H^{2}(\mathcal{M},\mathbb{R})}{\frac{2\pi}{q}H^{2}(\mathcal{M},\mathbb{Z})}\right|\det\hskip 0.0pt^{\prime}\left(\Delta_{1}^{T}\right)^{\frac{1}{2}}\left|\frac{H^{1}(\mathcal{M},\mathbb{R})}{\frac{2\pi}{q}H^{1}(\mathcal{M},\mathbb{Z})}\right|^{-1}\det\hskip 0.0pt^{\prime}\left(\Delta_{0}\right)^{-\frac{1}{2}}|U(1)|\;.$ (2.30)
@@
$\mathcal{Z}_{\rm PI}^{p}[\mathcal{M}]$ $=\sum_{\mathscr{F}\in\frac{2\pi}{q}\mathcal{H}^{p+1}(\mathcal{M},\mathbb{Z})}\hskip-14.22636pte^{-S[\mathscr{F}]}\int\frac{\mathcal{D}A}{|\mathcal{G}_{p-1}/(\mathcal{G}_{p-2}/\dots)|}\,e^{-S[dA]}$ (2.31)
$=\sum_{\mathscr{F}\in\frac{2\pi}{q}\mathcal{H}^{p+1}(\mathcal{M},\mathbb{Z})}\hskip-14.22636pte^{-S[\mathscr{F}]}\prod_{k=0}^{p}\left(\det\hskip 0.0pt^{\prime}\left(\Delta_{k}^{T}\right)^{-\frac{1}{2}}\left|\frac{H^{k}(\mathcal{M},\mathbb{R})}{\frac{2\pi}{q}H^{k}(\mathcal{M},\mathbb{Z})}\right|\right)^{(-)^{p-k}}\;.$
@@
$=|\mathcal{G}_{\rm nontriv}|\left|\frac{H^{p-1}(\mathcal{M},\mathbb{R})}{\frac{2\pi}{q}H^{p-1}(\mathcal{M},\mathbb{Z})}\right|\,\prod_{k=0}^{p-2}\left(\det\hskip 0.0pt^{\prime}\left(\Delta_{k}^{T}\right)^{-\frac{1}{2}}\left|\frac{H^{k}(\mathcal{M},\mathbb{R})}{\frac{2\pi}{q}H^{k}(\mathcal{M},\mathbb{Z})}\right|\right)^{(-)^{p-k+1}}$
where $\mathcal{G}_{\rm nontriv}$ is the nontrivial part of $\mathcal{G}_{p-1}$, i.e. $\frac{2\pi}{q}\mathcal{H}^{p}(\mathcal{M},\mathbb{Z})$ plus $d\lambda$ with $\lambda$ a non-closed (p1)(p-1)-form. Then the measure can be written simply as $\mathcal{D}A/|\mathcal{G}|$. For later use we evaluate $|U(1)|$ explicitly here. A natural measure for the scalar gauge parameter $\phi$ is $\mathcal{D}\phi=\prod_{n}\frac{\mu^{p+1}d\phi_{n}}{\sqrt{2\pi}}$, associated with the Laplacian eigendecomposition $\phi(x)=\sum_{n}\phi_{n}f_{n}(x)$. Here the arbitrary mass scale $\mu$ is analogous to a UV cutoff, and is included so that the measure is dimensionless. The normalized zero mode is $f_{0}(x)=V_{\mathcal{M}}^{-1/2}$, where $V_{\mathcal{M}}$ is the volume of $\mathcal{M}$. The zero mode coefficient then inherits the periodicity $\phi_{0}\sim\phi_{0}+\frac{2\pi}{q}V_{\mathcal{M}}^{1/2}$. We then have
$|U(1)|\equiv\int_{0}^{\frac{2\pi}{q}\sqrt{V_{\mathcal{M}}}}\frac{\mu^{p+1}}{\sqrt{2\pi}}d\phi_{0}=\frac{\mu^{p+1}}{q}\sqrt{2\pi V_{\mathcal{M}}}\;.$ (2.33)
@@
$\det\hskip 0.0pt^{\prime}\left(\Delta_{k}\right)=\det\hskip 0.0pt^{\prime}\left(\Delta_{k}^{T}\right)\det\hskip 0.0pt^{\prime}\left(\Delta_{k-1}^{T}\right)\;.$ (2.34)
The functional determinant factors tell us about the local degrees of freedom of the theory. We now carry out a sanity check to confirm that in total there are $D-2\choose p$} local degrees of freedom, as argued above. A $k$-form has $D\choose k$} independent components, so a determinant $\det^{\prime}\left(\Delta_{k}\right)^{-\frac{1}{2}}$ represents $D\choose k$} degrees of freedom. Similarly its reciprocal represents $D\choose k$} degrees of freedom. Let us write the number of degrees of freedom in a transverse $k$-form determinant as $n_{k}$. From (2.34) we apparently have $n_{k}+n_{k-1}={D\choose k}$. This recursion relation is solved by